Which of the following is equal to 5.93 x 10^-2

why the quiz i took says the right anwer is 0.0593 instead of 593.00 or just 593? why are the 0 in the left side? :S

Answer by Alan3354(6089)

Which of the following is equal to 5.93 x 10^-2
why the quiz i took says the right anwer is 0.0593 instead of 593.00 or just 593? why are the 0 in the left side? :S
----------------
10^-2 = 1/10^2 = 1/100 = 0.01 (different ways of writing it)
5.93 x 0.01 = 0.0593